OpenStudy (anonymous):

F=k*x^n where F is the force and x is the displacement find the value of n for the power done by the Force is independent of x?

OpenStudy (anonymous):

@electrokid @Vincent-Lyon.Fr

OpenStudy (anonymous):

P = F.V P = F. dx/dt ryt?

OpenStudy (anonymous):

right, so, you have \[P\propto\int F\mathrm{d}x\]

OpenStudy (anonymous):

so, I believe, when\[n=-1\], the power becomes independent of displacement!

OpenStudy (vincent-lyon.fr):

But \(\int F\mathrm{d}x\) is work done, not power. I do not understand what the problem is about.

OpenStudy (anonymous):

We need to find the value of n for which power by Force is independent of x

OpenStudy (anonymous):

x= displacement

OpenStudy (anonymous):

@Coolsector

OpenStudy (anonymous):

\[P=Fv=kx^nv\] \[\frac{\partial P}{\partial x}=0\] \[knx^{n-1}v=0\] \[n=0\] \[\ddot\smile\]